OpenStudy (anonymous):

12.Solve the equation 2x^2 + 10x = 0. Check your solution and state your final solution set. Help please this is the only question i seem to be stuck on!

OpenStudy (abb0t):

factor out an \(\sf \color{red}{x}\)

OpenStudy (campbell_st):

remove 2x as a common factor

OpenStudy (abb0t):

2x(x+5}=0 x = __ x = __

OpenStudy (anonymous):

so 2x^2+10x = 0 ?

OpenStudy (abb0t):

What? Lol No.

OpenStudy (anonymous):

im soo confused lol

OpenStudy (abb0t):

2x = 0 (x+5)=0 solve for \(\sf \color{red}{x}\)

OpenStudy (anonymous):

2x /2 = 0/ 2 x=0 (x+5) - 5 = 0-5 x= -5 ?

OpenStudy (abb0t):

Yes. Those are your solutions.

OpenStudy (anonymous):

thanks!!(: